# Orchid: Flexible and Data-Dependent Convolution for Sequence Modeling

### v2

@@ Abstract
In the rapidly evolving field of deep learning, the demand for models that are both expressive and computationally efficient has never been more critical. This paper introduces Orchid, a novel architecture designed to address the quadratic complexity of traditional attention mechanisms without compromising the ability to capture long-range dependencies and in-context learning. At the core of this architecture lies a new data-dependent global convolution layer, which contextually adapts its kernel conditioned on input sequence using a dedicated conditioning neural network. We design two simple conditioning networks that maintain shift equivariance in our data-dependent convolution operation. The dynamic nature of the proposed convolution kernel grants Orchid high expressivity while maintaining quasilinear scalability for long sequences. We evaluate the proposed model across multiple domains, including language modeling and image classification, to highlight its performance and generality. Our experiments demonstrate that this architecture not only outperforms traditional attention-based architectures such as BERT and Vision Transformers with smaller model sizes, but also extends the feasible sequence length beyond the limitations of the dense attention layers. This achievement represents a significant step towards more efficient and scalable deep learning models for sequence modeling.
@@ 1 Introduction
This paper proposes a novel data-dependent convolution mechanism to tackle the inherent quadratic complexity of traditional attention mechanisms, while maintaining the model’s ability to capture long-range dependencies and in-context learning. The data-dependent convolution layer contextually adapts its kernel based on input data using a dedicated conditioning neural network. We design two simple yet effective conditioning networks that maintain shift equivariance in the adaptive convolution operation. By combining these adaptive mechanisms with gating operations, our proposed model—named Orchid—achieves high expressivity while offering quasilinear scalability (with a complexity of $\mathcal{O}(L\log L)$) for long sequences. Evaluation across various domains, including language modeling and image classification, presented in section LABEL:sec:experiments and Appendix, demonstrates the Orchid architecture’s performance and generality, outperforming attention-based architectures, like BERT and Vision Transformers, with smaller model sizes. Moreover, its allows for handling very large sequence lengths that are beyond the limitations of the dense attention layers. This achievement lays the foundation for further advancements in more efficient and scalable sequence modeling architectures.
@@ 2 Background
Figure 2.1: Orchid block. $\odot$ and * denote element-wise multiplication and the convolution operator, respectively. The convolutions are implemented in the frequency domain using FFT. On the right two different conditioning networks, introduced in equations (2) and (3) as shift-invariant convolution kernels, are depicted. In this model, the convolution is performed efficiently in the spectral domain, so the kernel in the frequency domain, $h^{\mathcal{F}}=h_{0}^{\mathcal{F}}+h^{\mathcal{F}}_{\theta}({\bm{x}})$, is computed.
@@
Discrete-time linear convolution is a fundamental operation in digital signal processing that calculates the output as the weighted sum of the finite-length input ${\bm{x}}$ with shifted versions of the convolution kernel, ${\bm{h}}$, also known as the impulse response of a linear time-invariant (LTI) system.222 While generally ${\bm{x}}$ represents a sequence of D-dimensional embeddings with length $L$, all the sequence mixing operators, including data-dependent convolutions, Fourier transforms and depthwise 1D convolutions (Conv1d) are performed along the sequence dimension. Therefore, for clarity and without loss of generality, we assume $D=1$ and ${\bm{x}}~{}\in~{}\mathbb{R}^{L}$. For a full list of notation definition, please refer to Appendix LABEL:apdx:not_def. Formally, it can be written as
@@
Standard convolution layers, explicitly parameterized with a short kernel, struggle to capture long-range dependencies in sequential data. Extending the kernel to match the input length enables modeling such dependencies but leads to linear growth in parameter counts and quadratic computational complexity. To mitigate the parameter growth challenge, the kernel of global (a.k.a. long) convolution can be implicitly parameterized using a multilayer perceptron (MLP), a technique that has been shown to maintain sub-linear parameter scaling (karami2019invertible; romero2021ckconv; li2022makes; poli2023hyena). Furthermore, a key advantage of convolution operators is that, leveraging the convolution theorem, convolution operators can be efficiently computed in the frequency domain using Fast Fourier Transform (FFT) algorithms, thereby reducing the computational complexity to $\mathcal{O}(L\log L)$ (cooley1965FFT). Formally, the linear convolution can be expressed in the frequency domain as $\hat{{\bm{y}}}=\mathcal{F}^{-1}(\mathcal{F}(\hat{{\bm{h}}})\odot\mathcal{F}(\hat{{\bm{x}}}))={\bm{T}}^{-1}({\bm{h}}^{\mathcal{F}}\odot{\bm{T}}\hat{{\bm{x}}})$, where ${\bm{T}}$ is the DFT matrix, $\mathcal{F}$ denotes the discrete Fourier transformation, and $\hat{{\bm{x}}}$ denotes the zero-padded signal, defined as $\hat{{\bm{x}}}\triangleq\texttt{pad}_{2L}({\bm{x}})=[\mathbf{0}_{L};~{}{\bm{x}}]$. Additionally, the circular convolution can be simply computed as: ${\bm{y}}={\bm{h}}\circledast{\bm{x}}=\mathcal{F}^{-1}(\mathcal{F}({{\bm{h}}})\odot\mathcal{F}({{\bm{x}}}))$.
@@ 3.2 Preserving Shift-Equivariance in Data-Dependent Convolution
This property is particularly important because it ensures the operator’s response is robust to shift of features within the input, thereby enhancing the model’s generalization capabilities. This inductive bias is at the core of the widespread success of convolution operations (thomas1802tensor). Therefore, it is desirable to design conditioning network in the data-dependent convolution (1) to preserve shift equivariance property. To maintain this property for data-dependent convolution operations, it is sufficient to design filter kernel to be shift-invariant, i.e. $h(\texttt{shift}_{m}({\bm{x}}))=h({\bm{x}})$ (refer to Appendix LABEL:apdx:proofs for the proof). In the following, we present two conditioning network designs satisfying shift-invariance.
@@
[…]

### v3

@@ Abstract
In the rapidly evolving field of deep learning, the demand for models that are both expressive and computationally efficient has never been more critical. This paper introduces Orchid, a novel architecture designed to address the quadratic complexity of traditional attention mechanisms without compromising the ability to capture long-range dependencies and in-context learning. At the core of this architecture lies a new data-dependent global convolution layer, which contextually adapts its kernel conditioned on input sequence using a dedicated conditioning neural network. We design two simple conditioning networks that maintain shift equivariance in our data-dependent convolution operation. The dynamic nature of the proposed convolution kernel grants Orchid high expressivity while maintaining quasilinear scalability for long sequences. We evaluate the proposed model across multiple domains, including language modeling and image classification, to highlight its performance and generality. Our experiments demonstrate that this architecture not only outperforms traditional attention-based architectures such as BERT and Vision Transformers with smaller model sizes, but also extends the feasible sequence length beyond the limitations of the dense attention layers. This achievement represents a significant step towards more efficient and scalable deep learning models for sequence modeling. The code is available at https://github.com/Karami-m/orchid.
@@ 1 Introduction
This paper proposes a novel data-dependent convolution mechanism to tackle the inherent quadratic complexity of traditional attention mechanisms, while maintaining the model’s ability to capture long-range dependencies and in-context learning. The data-dependent convolution layer contextually adapts its kernel based on input data using a dedicated conditioning neural network. We design two simple yet effective conditioning networks that maintain shift equivariance in the adaptive convolution operation. By combining these adaptive mechanisms with gating operations, our proposed model—named Orchid—achieves high expressivity while offering quasilinear scalability (with a complexity of $\mathcal{O}(L\log L)$) for long sequences. Evaluation across various domains, including language modeling and image classification, presented in section 4 and Appendix, demonstrates the Orchid architecture’s performance and generality, outperforming attention-based architectures, like BERT and Vision Transformers, with smaller model sizes. Moreover, its allows for handling very large sequence lengths that are beyond the limitations of the dense attention layers. This achievement lays the foundation for further advancements in more efficient and scalable sequence modeling architectures.
@@ 2 Background
Figure 2.1: Orchid block architecture. This diagram illustrates the structure of the Orchid block. The core operation is a convolution (denoted by *), efficiently implemented in the frequency domain using FFT. Element-wise multiplication is denoted by $\odot$. On the right side, two different conditioning networks, introduced in equations (2) and (3) as shift-invariant convolution kernels, are illustrated. In this model, the convolution is performed efficiently in the spectral domain, so the kernel in the frequency domain, $h^{\mathcal{F}}=h_{0}^{\mathcal{F}}+h^{\mathcal{F}}_{\theta}({\bm{x}})$, is computed. The block also includes MLPs for linear projection and pointwise mixing of features at the beginning, that is common design choice used in various sequence modeling architectures.
@@
Discrete-time linear convolution is a fundamental operation in digital signal processing that calculates the output as the weighted sum of the finite-length input $\bm{x}$} with shifted versions of the convolution kernel, $\bm{h}$}, also known as the impulse response of a linear time-invariant (LTI) system. [^2] While generally $\bm{x}$} represents a sequence of D-dimensional embeddings with length $L$, all the sequence mixing operators, including data-dependent convolutions, Fourier transforms and depthwise 1D convolutions (Conv1d) are performed along the sequence dimension. Therefore, for clarity and without loss of generality, we assume $D=1$ and ${\bm{x}}~\in~\mathbb{R}^{L}$. For a full list of notation definition, please refer to Appendix A. Formally, it can be written as
@@
[…]
@@ 3.2 Preserving Shift-Equivariance in Data-Dependent Convolution
This property is particularly important because it ensures the operator’s response is robust to shift of features within the input, thereby enhancing the model’s generalization capabilities. This inductive bias is at the core of the widespread success of convolution operations (thomas1802tensor). Therefore, it is desirable to design conditioning network in the data-dependent convolution (1) to preserve shift equivariance property. To maintain this property for data-dependent convolution operations, it is sufficient to design filter kernel to be shift-invariant, i.e. $h(\texttt{shift}_{m}({\bm{x}}))=h({\bm{x}})$ (refer to Appendix B for the proof). In the following, we present two conditioning network designs satisfying shift-invariance.
@@
In our design, we deploy a hybrid spatial-frequency domain conditioning network. This network consists of a 1D depthwise linear convolution ($\texttt{Conv1d}()$)) with a short kernel length (typically 3-5) acting at the spatial domain, followed by a short convolution in the frequency domain. By operating in both spatial and frequency domains, the conditioning network effectively mixes information from neighboring tokens and spectral components. The resulting conditioning neural network is formulated as: